OpenStudy (anonymous):

Find the interval of increase and the interval of decrease of 200+8x^3+x^4

OpenStudy (xapproachesinfinity):

did you do first derivative?

OpenStudy (xapproachesinfinity):

\[(200+8x^3+x^4)'=24x^2+4x^3=4x^2(6+x)\]

OpenStudy (xapproachesinfinity):

set that equal zero for critical points

OpenStudy (xapproachesinfinity):

@Gaby7890 ?

OpenStudy (anonymous):

I forgot to put by using the second derivative test

OpenStudy (xapproachesinfinity):

second derivative shows concavity of the curve

OpenStudy (xapproachesinfinity):

here you just need interval of increase and decrease

OpenStudy (anonymous):

Yes, I need the intervals where it increase and the intervals where it decreases by using the second derivative test.

OpenStudy (xapproachesinfinity):

\(4x^2(6+x)=0 \Longrightarrow x=0 ~~or~~ x=-6\) critical points at x=0 and x=-6

OpenStudy (xapproachesinfinity):

now let's see \[\text{this expression } 4x^2(x+6) ~~\text{the sign here depends on} ~~(x+6)~~ \text{because }~~ 4x^2>0 ~~\text{for any x} \]

OpenStudy (anonymous):

Well my critical points are at X=0 and X=-4 by finding the Second derivative of 200+8x^3+x^4

OpenStudy (xapproachesinfinity):

hmm how come you get x=-4? can you show the work

OpenStudy (xapproachesinfinity):

as i said there no need for second derivative test second derivative test is used to show how the graph is curving concave up or down and show inflection point

OpenStudy (xapproachesinfinity):

you don't need second derivative

OpenStudy (anonymous):

f(X)=200+8x^3+x^4 f'(X)=24x^2+4x^3 f''(X)=48x^2+12x^2 =12x(X+4)

OpenStudy (xapproachesinfinity):

so now we know that f'(x)<0 fo x<-6 and f(x)>0 for x>-6

OpenStudy (xapproachesinfinity):

correction f'(x)>0 for x>-6

OpenStudy (anonymous):

X=0 and X=-4 are my inflection points

OpenStudy (xapproachesinfinity):

x=-4 is the where the inflection point happens

OpenStudy (xapproachesinfinity):

yes that's true

OpenStudy (xapproachesinfinity):

but your question is asking interval of increase and decrease not inflection point

OpenStudy (anonymous):

I'm trying to find the the intervals where it increases

OpenStudy (xapproachesinfinity):

you are doing an unnecessary work

OpenStudy (xapproachesinfinity):

i just did it we have f'<0 for x<-6 f'>0 for x>-6 meaning f is increasing for x>-6 and decreasing for x<-6

OpenStudy (xapproachesinfinity):

here is the graph that shows same result i got http://www.wolframalpha.com/share/clip?f=d41d8cd98f00b204e9800998ecf8427e8jc7pjgfe5

OpenStudy (anonymous):

Okay thank you for the help. By any chance do you know how to do 4csc(3x/2) with the intervals (0,2pi)?

OpenStudy (xapproachesinfinity):

as you can see at x=0 and at x=-4 inflection points as you said exactly

OpenStudy (xapproachesinfinity):

same principle

OpenStudy (anonymous):

Okay but this is another problem I need help with.

OpenStudy (xapproachesinfinity):

\[(4\csc(3x/2))'=-6\csc(3x/2)\tan(3x/2)\]

OpenStudy (xapproachesinfinity):

now you just look at the sign of tan(3x/2) for x in [0,2pi) tan is positive for [0,pi/2) and [pi, 3pi/2) for csc which is just 1/sin positive in first quadrant and second quadrant and negative in III, IV

OpenStudy (xapproachesinfinity):

you just find then the sign of the product

OpenStudy (anonymous):

Would trying to find the second derivative of 4csc(3x/2) change anything. Because I have to find the (leftmost, interval, smaller X value) and also I have to find the (leftmost interval, larger X value)?

OpenStudy (anonymous):

By using the second derivative test.

OpenStudy (xapproachesinfinity):

hmm i don't know what you mean by leftmost? any example

OpenStudy (xapproachesinfinity):

oh correction derivative should have cotangent not tangent sorry blind it there i was thinking somehow of sec instead of csc

OpenStudy (xapproachesinfinity):

should be -6cot(3x/2)csc(3x/2) correction but the my talk about signs still holds since cot is just 1/tan

OpenStudy (anonymous):

Consider the function below on the interval (0, 2π). (Give your answers correct to 2 decimal places.) f(x) = 4 csc((3x)/2) (a) Find the points of inflection of the function. (If none exist, enter NONE.) (b) Find the open intervals where the graph is concave upward. ( Correct: Your answer is correct. , Incorrect: Your answer is incorrect. ) (leftmost interval, smaller x value) ( , Correct: Your answer is correct. ) (rightmost interval, larger x value) (c) Find the open interval where the graph is concave downward. ( Incorrect: Your answer is incorrect. , Incorrect: Your answer is incorrect. )

OpenStudy (xapproachesinfinity):

oh yeah you definitely need second derivative for this one it is asking for inflection points those leftmost and rightmost they are just asking where it is concaved up that happens in two interval here that's why they give you two and then last question where it is concaved down

OpenStudy (xapproachesinfinity):

here we don't need first derivative to answer this we just do it and skip it and derive second derivative to answer this problem

OpenStudy (xapproachesinfinity):

it needs some more work to clean! i gotta go anyway to sleep

OpenStudy (xapproachesinfinity):

good luck weetie!

OpenStudy (anonymous):

Okay thank you!

OpenStudy (xapproachesinfinity):

no problem